OpenStudy (anonymous):

3x - 2 x + 2 = 3x x - 1

OpenStudy (mertsj):

\[\frac{3x-2}{x+2}=\frac{3x}{x-1}\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Yes.

hero (hero):

Okay, so @shylagill, what would be the next step?

OpenStudy (anonymous):

Okay, so @hero that's what I'm trying to figure out!

hero (hero):

Ever heard of cross multiplication?

OpenStudy (anonymous):

You cross multiply. :) So it goes:  Sorry, my writing is really bad.. :D Then, you solve the equation..